The following tables shows the mpg and curbed weight, in thousands for eight randomly selected vehicles:

MPG: 28 22 20 34 41 15 19 15

Weight: 2730 2860 3140 2180 2310 4450 3920 2590

Using XLS, calculate the correlation coefficient for this sample. Using a significance of .1, test the significance of the population correlation coefficient is negatively related between mpg and curbed weight. What conclusions can you draw?
Answer #1
MPG (X) Weight (Y) X * Y X2 Y2
28 2730 76440 784 7452900
22 2860 62920 484 8179600
20 3140 62800 400 9859600
34 2180 74120 1156 4752400
41 2310 94710 1681 5336100
15 4450 66750 225 19802500
19 3920 74480 361 15366400
15 2590 38850 225 6708100
Total 194 24180 551070 5316 77457600

r = ( ( N * \Sigma XY) - \Sigma X \Sigma Y) / (\sqrt{ (N \Sigma (X^{2}) - (\Sigma X)^{2} )(N \Sigma (Y^{2}) - (\Sigma Y)^{2} })
r = ( ( 8 * 551070 ) - 194 * 24180 ) / ( \sqrt{ ( 8 * 5316 ) - ( 194 )^{2} )( 8 * 77457600 ) - ( 24180 )^{2} })
Corelation Coefficient r = - 0.6825

To Test :-

H0 :- ρ = 0
H1 :- ρ < 0

Test Statistic :-
t = (r * √(n - 2) / (√(1 - r2))
t = ( -0.6825 * √(8 - 2) ) / (√(1 - 0.4658) )
t = -2.2873


Test Criteria :-
Reject null hypothesis if t < -t(α, n - 2 )
t(α, n-2) = t(0.1 , 8 - 2 ) = 1.4398
t < -t(α,n-2) = -2.2873 < -1.4398
Result :- Reject null hypothesis


Decision based on P value
P - value = P ( t > 2.2873 ) = 0.0311
Reject null hypothesis if P value < α = 0.1 level of significance
P - value = 0.0311 < 0.1 ,hence we reject null hypothesis
Conclusion :- Reject null hypothesis

There is sufficietn evidence to conclude that negatively related between mpg and curbed weight.